OpenStudy (anonymous):

Using the distance formula whats the distance to the nearest tenth between A(-7,-4) and b(-2,0)

OpenStudy (anonymous):

distance formula is \[\sqrt{(x-x)^2+(y-y)^2}\] \[\sqrt{(-7+2)^2+(-4-0)^2}\] \[\sqrt{(-5)^2+(-4)^2}\] \[\sqrt{(25)+(16)}\] \[\sqrt{31}\]

OpenStudy (anonymous):

oops \[\sqrt{41}\]

hero (hero):

Distance formula sucks

OpenStudy (anonymous):

\[\sqrt{41}\approx6.4\]